The sum of 3 integers is 251. The sum of the 1st and 2nd integers exceeds the 3rd by 45. The 3rd integer is 42 less than the 1st. Find the 3 integers
Answer by richard1234(7193)

We have the system

a + b + c = 251
a + b = c + 45
c = a + 42

Substituting the second equation into the first, we obtain

(c + 45) + c = 251 --> 2c + 45 = 251 --> c = 103. Replacing c into the third equation we get a = 61, and it follows that b = 87.